3. A company wanted to know if attending a course on how to be a successful salesperson can increase the average sales of itsPlease show work :)
Answer #1

Solution:

Here, we have to use paired t test.

The null and alternative hypotheses for this test are given as below:

H0: µd = 0 versus Ha: µd > 0

This is a right tailed test.

We take difference as after minus before.

Test statistic for paired t test is given as below:

t = (Dbar - µd)/[Sd/sqrt(n)]

From given data, we have

µd = 0

Dbar = 4.1667

Sd = 2.6394

n = 6

t = (Dbar - µd)/[Sd/sqrt(n)]

t = (4.1667 - 0)/( 2.6394/sqrt(6))

t = 3.8668

df = n – 1 = 5

α = 0.01

The p-value by using t-table is given as below:

P-value = 0.0059

P-value < α = 0.01

So, we reject the null hypothesis

There is sufficient evidence to conclude that the mean weekly sales for all salespersons increase as a result of attending this course.